Which of the following is the SOA (strongest oxidizing agent) according to the redox table? 


  1. Cu⁺
  2. Cu
  3. Cu²⁺
  4. Cu³⁺




1
Expert's answer
2022-01-10T05:51:03-0500

Solution:

1) Cu+(aq) + e → Cu(s), Eo = +0.52 V

2) Cu(s) - 2e → Cu2+(aq), Eo = -0.34 V

3) Cu2+(aq) + 2e → Cu(s), Eo = +0.34 V

4) Cu3+(aq) + 2e → Cu+(aq), Eo = +1.28 V


Cu3+ is the most effective oxidizer, having the largest positive electrode potential. 


Answer: Cu3+ is the strongest oxidizing agent.